OpenStudy (anonymous):

It takes 8 minutes for Byron to fill the kiddie pool in the backyard using only a handheld hose. When his younger sister is impatient, Byron also uses the lawn sprinkler to add water to the pool so it is filled more quickly. If the hose and sprinkler are used together, it takes 5 minutes to fill the pool. Which equation can be used to determine r, the rate in parts per minute, at which the lawn sprinkler would fill the pool if used alone?

OpenStudy (anonymous):

Pease help... Will give medal

OpenStudy (anonymous):

5/8 + 5r = 8 5/8+ 5r = 1 5(5/8)5() = r 5/8 = 5r

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@DaverB

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

PLEASE

OpenStudy (welshfella):

1/8 + 1/r = 1/5